# MalAlgoQA: A Pedagogical Approach for Evaluating Counterfactual Reasoning Abilities

### v1

@@ 5 Discussions
Table 7 analyzes the MIA performance of GPT-4o and LLaMA3-70B models across different grade levels in math and reading datasets. Both models consistently show a decline in performance with increasing grade levels, with LLaMA3-70B exhibiting some unexpected variability, particularly excelling in Grade 8 of the reading dataset. These results highlight the increasing challenges in tackling malgorithm identification task as the questions becomes more complicated.
@@
Figure 2: An illustration of MIA performance for LLMs across five math content classifications (Number & Operation, Algebra, Geometry & Measurement, Data Analysis, Probability) and two reading content classifications (Informational Text, Literature).
6 Related Works
Counterfactual reasoning is a cognitive process that involves considering alternative scenarios to understand causality, which has gained increasing attentions with the advent of artificial intelligence. Pearl’s foundational work on causal inference [12] has established a theoretical framework for understanding counterfactuals, which has been applied to various domains, including fairness, interpretability, and robustness in machine learning [8, 17, 24].
In the field of NLP, counterfactual reasoning is crucial for tasks that require understanding hypothetical scenarios and the causal relationships between events. However, most existing NLP models focus on predictive tasks rather than for hypothetical reasoning [3, 2, 6]. While several datasets have been created to evaluate different aspects of reasoning in language models, relatively few focus explicitly on counterfactual reasoning. For instance, the bAbI dataset [28] includes tasks that require basic forms of logical reasoning, while the SQuAD dataset [15] focuses on reading comprehension. However, these datasets do not emphasize counterfactual scenarios. Recent efforts like the COPA dataset [16] and the CQA dataset [22] have introduced questions that require commonsense reasoning, which often involves understanding causal relationships. Nonetheless, these datasets are not explicitly designed to test counterfactual reasoning. To address this gap, we introduce MalAlgoQA, a novel dataset specifically designed to evaluate the counterfactual causal reasoning capabilities of LLMs through a pedagogical approach with multiple-choice question answering format.
@@ 6 Related Works
Applications of MalAlgoQA for personalized education: The use of LLMs in classrooms [20, 21], for instance, could profoundly influence the learning experiences of students. While LLMs can provide personalized learning support and instant feedback, their inability to accurately identify and address student misconceptions, as highlighted by our study, could potentially reinforce incorrect understanding or reasoning. However, the transparency of LLMs’ decision-making process is a crucial ethical consideration. If LLMs are to be trusted educational aids, it is critical for students, teachers, and parents to understand the reasoning behind the LLMs’ responses. Our work with the MalAlgoQA dataset is a significant step towards understanding and addressing these ethical considerations. By examining the counterfactual causal reasoning abilities of LLMs, we can assess their readiness for use in educational settings and contribute to the ongoing discourse surrounding the ethical deployment of AI in education. However, we acknowledge that this is a complex, multifaceted issue that requires continuous exploration, vigilance, and open dialogue among researchers, educators, policymakers, and society at large.
7 Conclusion and Future Works
This paper presents MalAlgoQA, a novel dataset designed to critically evaluate the counterfactual reasoning capabilities of LLMs through a pedagogical prespective. Additionally, we introduce the Malgorithm Identification task along with associated metrics for its evaluation. Our findings underscore the need for ongoing research and development the ability to identify malgorithms, particularly for educational applications where recognizing flawed reasoning processes is essential. Notably, our study also challenges the efficacy of the chain-of-thought prompting technique in enhancing counterfactual causal reasoning, advocating for further investigation into suitable prompting techniques for robust malgorithm identification. In the future, we plan to expand the MalAlgoQA dataset to cover a broader range of subjects and reasoning complexities. We also aim to explore different training methodologies and architectural modifications to LLMs that may enhance their ability to engage in counterfactual causal reasoning, ultimately contributing to the development of more cognitively-aligned language models.
@@ Acknowledgments
This work was supported by NSF grants 1842378, ONR grant N0014-20-1-2534, AFOSR grant FA9550-22-1-0060, and a Vannevar Bush Faculty Fellowship, ONR grant N00014-18-1-2047.
Appendix A Additional Analysis
A.1 Grade vs DOK for different content classifications
Figure 3: An image of MIA performance across different Grades and DOK levels for various Content Classifications for Math question sets. The top number represents MIA performance and the bottom number represents question counts. The data analysis and probability classifications are combined together for illustration. (best viewed in colors).
Figure 4: An image of MIA performance across different Grades and DOK levels for two Content Classifications for Reading question sets. The top number represents MIA performance and the bottom number represents question counts (best viewed in colors).
Figures 3 and 4 illustrate the relationships between grade levels, DOK for different content classifications in math and reading question sets. In mathematics, Number & Operation and Algebra questions demonstrate relatively consistent performance, with higher DOK levels often showing better results. For instance, MIA performance for Number & Operation peaks in lower grades and declines in higher grades, particularly Grade 7. This trend suggests increasing complexity in numerical operations as grade levels increases. On the other hand, Geometry & Measurements, Data Analysis and Probability exhibit high variability, with strong MIA performance in some grades but gaps in others. In reading, performance for Informational Text is generally consistent across grades, peaking in Grade 8, with lower DOK levels typically yielding higher MIA performance. Literature shows high variability, reflecting the varying complexity of literary analysis tasks. Notably, content classifications in mathematics show more consistent patterns across grades compared to reading ones.
A.2 Performance across Grade level and Content Classifications
Figure 5: An illustration of MIA performance across different Grades and Content Classifications for Math and Reading question sets. The top number represents MIA performance and the bottom number represents question counts (best viewed in colors).
Figure 5 shows the relationship between grade levels and content classifications for math and reading question sets. In mathematics question set, Algebra demonstrates relatively consistent performance across all grades, with the number of algebra questions increasing significantly in higher grades, peaking at 191 data points in Grade 8. Number & Operation exhibits a gradual decline in MIA performance from lower to higher grades. In contrast, Data Analysis and Probability, Geometry & Measurement classifications show varied MIA performance across grade levels. For instance, Data Analysis and Probability demonstrates exceptionally high performance in Grade 6 (81.82), albeit with relatively few data points (22). In reading question set, MIA performance for both Informational Text and Literature classifications fluctuates across grades without much correlation. Informational Text achieves a significant peak in MIA performance for Grade 8, while Literature questions in Grade 5 demonstrate exceptionally high performance. This variability could be attributed to differences in text complexity or question difficulty at different grade levels for the Malgorithm Identification problem.
A.3 Performance across Questions
Figure 6: The distribution of GPT4-o performance in identifying malgorithms across question levels for math and reading question sets. The x-axis represents the number of correctly identified malgorithms per question, ranging from 0 to 3.
We can find the performance of LLMs in identifying malgorithms across each question in both mathematics and reading comprehension sets in Figure 6. The number of questions increases as we progress from 0 to 3 correctly identified malgorithms per question. This pattern suggests that LLMs are more likely to identify multiple malgorithms correctly rather than missing all or most of them. However, despite this positive trend, it is crucial to note that LLMs’ overall performance on this task remains suboptimal. The presence of the significant number of questions at the 0 and 1 levels across both domains indicates room for improvement. In mathematics, the combined number of questions with 0 or 1 correctly identified malgorithms is comparable to those with 3 correct identifications. Similarly, in reading comprehension, while there is a clearer upward trend, the number of questions with 0 or 1 correct identifications remains considerable. These findings underscore the need for further research of LLMs’ counterfactual reasoning capabilities in malgorithm identification across various question types and subject areas.
Appendix B Examples of Mathematics dataset
We provide five examples, one per content classifications, of question, choices and their corresponding rationales in Table 8 and 9.
@@ Appendix B Examples of Mathematics dataset
Appendix C Examples of Reading Comprehension dataset
We provide two examples of passage, question, choices and their corresponding rationales in Table 10 and 11.
Reading Question Set Example 1:
Content Classification: Literature
Passage: Blueberries
1 You ought to have seen what I saw on my way
2 To the village, through Pattersons pasture to-day:
3 Blueberries as big as the end of your thumb,
4 Real sky-blue, and heavy, and ready to drum
5 In the cavernous pail of the first one to come!
6 And all ripe together, not some of them green
7 And some of them ripe! You ought to have seen!
Question: In lines 1-7, the speaker expresses great surprise about the ...?
Answer and Rationales:
A: large pails of blueberries.
Rationale A: Some readers might find this option plausible because the speaker mentions the cavernous pail of the first one to come" (line 5). However, this mention of the pail is referring to the pail into which the lucky picker who first stumbles across this crop will put the Blueberries as big as the end of your thumb" (line 3). The speaker does not express surprise that pickers are bringing large pails.
B: number of blueberry pickers.
Rationale B: Some readers might choose this option because the phrase the first one to come" in line 5 refers to a blueberry picker. However, although the speaker is predicting success for anyone who comes to pick the blueberries, the speaker is not expressing surprise that there are no pickers yet. Based on the speakers description of how big, blue, and ripe the berries are (as big as the end of your thumb, Real sky-blue, and heavy, and ready" (lines 3 and 4)), the berries are undisturbed, and no one has discovered them except for the speaker.
C: ripeness of the blueberries.
Rationale C: The speaker describes big as the end of your thumb" (line 3) blueberries that are Real sky-blue, and heavy," (line 4) and all ripe together" (line 6). The speaker repeats, You ought to have seen" in lines 1 and 7 to convey surprise.
D: size of new blueberry bushes.
Rationale D: This option might seem plausible because the size of the blueberries and their bushes are described. The first speaker draws attention to size when describing the blueberries as big as the end of your thumb" (line 3). In line 13, the second speaker exclaims, Why, there hasnt been time for the bushes to grow." This line implies that the bushes must be small, so it is surprising that the blueberries they produce are numerous and as big as the end of your thumb." Although lines 1-7 address the blueberries, the speaker does not mention the bushes on which they grow in those lines.
Table 10: Example 1 of Reading question set.
Reading Question Set Example 2:
Content Classification: Informational Text
Passage: Creating Codes: Grace Hopper and the Computer Compiler
1 The computer screen freezes. No matter what or how often you click, the hourglass or rotating circle (that symbol that insists, Im working on it!") spins away. You want to shake the screen. Someone next to you says, I think you found a bug."
Question: What is the main purpose of paragraph 1?
Answer and Rationales:
A: To demonstrate the importance of the subject
Rationale A: Although some readers may conclude that the frozen computer screen in paragraph 1 demonstrates how important computer programming is currently, the paragraph describes a familiar feeling of frustration rather than detailing the more serious consequences that could result from programming mistakes.
B: To help readers make a connection with the topic
Rationale B: The paragraph describes a person facing a frozen computer screen that many readers have probably experienced in order to help readers make a personal connection to the topic. By describing this common situation, the readers interest is engaged. So, when the article continues on to discuss Hoppers contribution to programming and how she helped solve some of the early computer bugs, readers feel more connected with the topic.
C: To give background information about the subject
Rationale C: Although introductions often provide background information, and some readers may conclude that explaining a computer bug is background information for the article, the topic of the article focuses not on computer bugs but on Hoppers contributions to computer programming.
D: To provide readers with a brief summary of the topic
Rationale D: Although the main topic of the article is Hoppers contribution to computer programming and paragraph 1 mentions computers, paragraph 1 describes a common situation in order to raise interest among readers. It describes a person having computer problems rather than summarizing Hoppers contributions to computer programming.
Table 11: Example 2 of Reading question set

### v2

@@ 4.2 Main Findings
Ineffectiveness of Few-Shot Learning
The data clearly demonstrates that few-shot learning does not significantly improve MIA performance across models. For GPT-4o on the Math dataset, the MIA score with CoT prompting actually decreases slightly from $66.10\%$ (zero-shot) to $65.10\%$ (5-shot). LLaMA3-70B shows a similar trend, with MIA dropping from $55.32\%$ to $50.6\%$. Even in cases where there is a marginal improvement, the gains are negligible. For instance, LLaMA3-8B’s MIA increases from $27.05\%$ to $31.39\%$ with 5-shot CoT prompting - a difference of only $4.34$ percentage points. GPT-3.5 shows a similarly small improvement from $14.74\%$ to $17.06\%$.
These results suggest that providing examples does not substantially enhance an LLM’s ability to identify malgorithms. This finding is particularly noteworthy given that few-shot learning often improves performance on other NLP tasks. The ineffectiveness of this approach for malgorithm identification underscores the unique challenges posed by this task and suggests that more sophisticated techniques may be needed to improve LLM performance in this area.
4.3 Understanding CoT Prompting for MIA
Our analysis of Chain-of-Thought (CoT) prompting reveals an unexpected ineffectiveness in improving Malgorithm Identification Accuracy (MIA), and in some cases, even leads to decreased performance compared to simple prompting.
Inconsistent impact across subjects: For GPT-4o, CoT prompting shows negligible difference in Math MIA ($65.1\%$ vs $65.0\%$ for simple prompting), but significantly underperforms in Reading MIA ($68.6\%$ vs $82.8\%$ for simple prompting).
Consistent underperformance of LLaMA-70B: LLaMA-70B demonstrates lower MIA prompting in both Reading ($58.4\%$ vs $60.0\%$) and Math ($48.3\%$ vs $50.6\%$) with CoT compared to simple prompting.
Minimal impact on AIA: CoT prompting generally performs similarly to simple prompting for Algorithm Identification Accuracy (AIA) across models and subjects, indicating its ineffectiveness is specific to malgorithm identification.
Figure 2: An illustration of MIA performance across five math content classes (Number & Operation, Algebra, Geometry & Measurement, Data Analysis, Probability) and two Reading content classes (Informational, Literature).
These results challenge the assumption that explicit step-by-step reasoning always improves performance on complex tasks. The data suggests that CoT prompting, while effective for many NLP tasks, may actually hinder the model’s ability to identify flawed reasoning paths in the context of malgorithm identification.
A plausible explanation for this counterintuitive finding lies in the nature of the training data typically used for LLMs. These models are predominantly exposed to examples of correct reasoning during pre-training and fine-tuning. Consequently, when CoT prompting guides the model through a structured thought process, it likely leverages these learned patterns of correct reasoning. This may inadvertently anchor the model to correct logical pathways, making it more challenging to deviate from these patterns and identify flawed reasoning. In essence, CoT prompting might be reinforcing the model’s bias towards correct reasoning, which is counterproductive when the task specifically requires identifying incorrect logical steps.
5 Analysis: Educational Dimensions
@@
Table 7 analyzes the MIA performance of GPT-4o and LLaMA3-70B models across different grade levels in math and reading datasets. Both models consistently show a decline in performance with increasing grade levels, with LLaMA3-70B exhibiting some unexpected variability, particularly excelling in Grade 8 of the reading dataset. These results highlight the increasing challenges in tackling malgorithm identification task as the questions becomes more complex.
@@
6 Implications for AI in Education
The findings from our study on malgorithm identification have significant implications for the development and application of AI in educational contexts:
Limitations in Understanding Student Misconceptions: The substantial performance gap between Algorithm Identification Accuracy (AIA) and Malgorithm Identification Accuracy (MIA) suggests that current LLMs are far more adept at recognizing correct reasoning than identifying flawed logic. For GPT-4o, we observed a drop from $95.65\%$ AIA to $66.10\%$ MIA on the Math dataset. This disparity indicates that AI tutoring systems based on these models may excel at confirming correct answers but struggle to diagnose and address student misconceptions effectively. This limitation could lead to missed opportunities for targeted intervention and personalized learning.
Rethinking Training Approaches for Error Recognition The ineffectiveness of CoT prompting in improving MIA performance reveals a fundamental challenge in training LLMs to recognize and analyze errors. Rather than simply diversifying training data, this finding suggests a need for novel training paradigms that specifically target the skill of error identification Sonkar et al. (2024a). Future AIED systems may require specialized fine-tuning processes or architectural modifications that enable models to simultaneously reason about correct and incorrect logical pathways. This could involve developing new pre-training tasks focused on error detection, or creating model architectures that maintain separate representations for correct and flawed reasoning patterns. Such innovations could lead to AI tutors that are equally adept at reinforcing correct understanding and diagnosing misconceptions, significantly enhancing their effectiveness in personalized education.
Rethinking Feedback Mechanisms: The challenges in malgorithm identification suggest that current AI models might provide incomplete or misleading feedback to students. AIED systems need to be designed with an awareness of these limitations, potentially incorporating human oversight or alternative assessment methods to ensure that student misconceptions are accurately identified and addressed.
Implications for Automated Grading and Assessment: The performance discrepancy between AIA and MIA also raises concerns about the reliability of AI-driven automated grading systems, particularly for open-ended questions or problems requiring complex reasoning. Educational institutions considering the implementation of such systems should be aware of these limitations and consider alternate Sonkar et al. (2024b, d) or hybrid approaches with human review.
Adaptive Difficulty Handling: The observed decline in MIA performance across increasing grade levels and DOK levels highlights the need for AIED systems that can adapt to varying levels of question complexity. Future AI tutors should be designed with mechanisms to maintain consistent performance across different difficulty levels, ensuring effective support for students as they progress through more challenging material.
The difficulty LLMs face in identifying flawed reasoning points to a broader challenge in counterfactual reasoning. For AIED applications, this suggests a need to develop models with enhanced capabilities in exploring alternative scenarios and reasoning paths. Such improvements could lead to more sophisticated tutoring systems that can not only identify errors but also explain why certain approaches are incorrect and guide students towards correct understanding. By addressing these implications, AIED researchers can work towards not only improving the accuracy of student assessment but also enhancing the ability of AI systems to provide targeted and personalized support.
7 Related Work
Counterfactual reasoning, a critical cognitive process for understanding causality, has gained significant attention in machine learning (ML) and NLP research. Pearl’s seminal work on causal inference Pearl (2009) has been widely applied to address challenges in fairness Kusner et al. (2017), interpretability Wachter et al. (2018), and robustness Bareinboim and Pearl (2016) in ML. In NLP, counterfactual reasoning is important for tasks requiring comprehension of hypothetical scenarios. However, most existing NLP models primarily focus on predictive tasks rather than hypothetical reasoning Bommasani et al. (2021); Bender et al. (2021). Recent work has begun to address this gap developing specialized datasets and models for counterfactual reasoning Qin et al. (2019). Several datasets have been created to evaluate different aspects of reasoning, including counterfactual reasoning such as COPA dataset Roemmele et al. (2011), Counterfactual Story Rewriting task Qin et al. (2020), and CQA dataset Talmor et al. (2019). These efforts are complemented by development of new modeling approaches, such as causal LMs Feder et al. (2021), counterfactual data augmentation Kaushik et al. (2020), and generating counterfactual explanations Wu et al. (2021); Ross et al. (2021). Despite these advancements, challenges remain such as handling multi-step causal chains and generalizing across diverse domains Xu et al. (2023); Keith et al. (2020); Feder et al. (2022).
8 Conclusion and Future Work
This paper introduces MalAlgoQA, a novel dataset designed to evaluate the counterfactual reasoning capabilities of LLMs through a pedagogical lens. By proposing the Malgorithm Identification task and its associated metrics, we provide a new framework for assessing LLMs’ ability to recognize and analyze flawed reasoning patterns. Our findings reveal significant challenges in counterfactual reasoning for state-of-the-art LLMs. These results highlight the need for further research into enhancing LLMs’ counterfactual reasoning abilities, particularly for educational applications where understanding student misconceptions is crucial. Future work will focus on expanding the dataset, exploring novel training methodologies, and investigating architectural modifications to develop more cognitively-aligned language models capable of robust counterfactual reasoning.
9 Limitations
While our research sheds light on the counterfactual reasoning capabilities of LLMs, we acknowledge the inherent complexity and intricacy of the cognitive process may not be fully captured by our current evaluation framework. The MalAlgoQA dataset and the Malgorithm Identification task, although designed with careful consideration, are initial steps toward understanding the intricate aspect of cognition. The chain-of-thought (CoT) prompting technique is not a panacea and its effectiveness vary depending on the specific context and task. As research in this field progresses, we anticipate further refinements and enhancements to our methodology, contributing to a more comprehensive understanding of counterfactual reasoning in LLMs.
10 Ethics and Risk
@@
The use of LLMs in classrooms, for instance, could profoundly influence the learning experiences of students. While LLMs can provide personalized learning support and instant feedback, their inability to accurately identify and address student misconceptions, as highlighted by our study, could potentially reinforce incorrect understanding or reasoning. However, the transparency of LLMs’ decision-making process is a crucial ethical consideration. If LLMs are to be trusted educational aids, it is critical for students, teachers, and parents to understand the reasoning behind the LLMs’ responses. Our work with the MalAlgoQA dataset is a significant step towards understanding and addressing these ethical considerations. By examining the counterfactual causal reasoning abilities of LLMs, we can assess their readiness for use in educational settings and contribute to the ongoing discourse surrounding the ethical deployment of AI in education. However, we acknowledge that this is a complex, multifaceted issue that requires continuous exploration, vigilance, and open dialogue among researchers, educators, policymakers, and society at large.
@@ Acknowledgments
This work was supported by NSF grant 1842378, ONR grant N0014-20-1-2534, AFOSR grant FA9550-22-1-0060, a Vannevar Bush Faculty Fellowship, and ONR grant N00014-18-1-2047.
Appendix A Appendix
Examples of MalAlgoQA dataset
We provide five math examples, one per content classifications, of question, choices and their corresponding rationales in Table 8 and 9.